Shift summary Evening 30 Nov. 1999

Nickolas Solomey (solomey@hep.uchicago.edu)
Wed, 1 Dec 1999 00:01:57 -0600

Shift Summary for evening shift, 30 Nov. 1999

On Shift: Rick T. and Nick

Hours of beam delivered at all intensities: 4.25
Hours of beam delivered at nominal intensity: 3.5
Hours of E799 data (special runs not included): 3.5
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Tevatron accesses, plus Trigger problems in
DAQ, NM2Q1 480 V wall power tripped,

========================================================================
Runs ended this shift: 15096,15097,15098,15099,
15100 is a long good run started this shift but on-going.
Tapes brought to FCC this shift: none
(including tapes for online split)
========================================================================
Accesses during shift: Sasha, Rick T., Valeri
========================================================================
Overview of shift: Debugging Trigger problem while waiting for
Tevatron accesses to finish.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15096 15:30 New shift takes over, but we had both been
here for a while before the start of our shift.
133 Some small amount of beam starts to arrive, but
the old one was a cosmic run, so we start a
new run.
15097 15:40 Beam back trying to start a new run, having
problems with trigger, idel L2 busy.
16:00-:45 Elmhurst College students came by for a tour,
no access just control room.
15098 17:40 Still working on debugging trigger problems.
18:00 Tevatron down preparing for access to
disconect A0 straight section short.
15099 18:15 Asking for accesses to work on VV' latches.
18:55 Trigger problem is now gone, but it was
solved by wiggling cables and resetting them
so it is suspicious.
18:59 Tevatron still down with access.
19:15 MCR called, tevatron access almost over they
asked us if we could end of access; Valerie
and Sasha go back down to undo their VV'
latches uncabling.
19:35 MCR called, saying they are ready to start
Tevatron.
19:36 Channel 13 new message says 45 more min till
beam
20:15 We come out of access mode, Tevatron beam
expected soon.
15100 6 20:30 First time tonight DA ready and beam from
Tevatron at 8.2e12, spill structure is
reasonable.
14 Tevatron missed a spill.
15 This spill beam aborted half way threw.
16&17 No beam.
18 Beam Back.
64 21:50 No beam to KTeV, our magnets NM2Q1 and NM2Q2
tripped off, MCR is coming out to look at it.
22:00 480 Volts wall unit tripped, electricans are
coming in to reset, afterwhich they will find
out if anything serious caused it.
22:10 480 V reset and NM2Q1 and NM2Q2 came back on
normally with no sign of a problem.
82 22:14 Beam is back.
120 Problems noticed on historgram scan:
1)V latch in time slice is drifting.
2)Sigma for SOD in DC3X and DC4Y are out of
limits.
3)CsI channel 1497 has kind of large RMS.
125 Tevatron missd a spill.
126 Noticed online plot of DC4 has a funny looking
timing distrubtion (possibly RF pickup)
with a ripply structure, looks worse than DC1
which is not a good indication.
156 Last 4 spills lots of out of limit alarms.
Tevatron had given us spills with 1.07e13 and
1.13e13, we called and asked to make it less
than 1.00e13.
159 No beam that spill.
161 1.06e13!
162 End of our shift, beam is hot and tapes almost full.

========================================================================
NOTE to owl shift from Sasha (run manager): "don't do muon run or flip
magnet tonight, just keep data taking." Valeri said that start of next run
should ask you to update adgn constants answer yes, if it does not ask
force it.
========================================================================
NOTE to tomorrow's day shift:

Rick C. phoned saying that KAMI beam test for tomarrow is canceled,
KTeV should continue with HEP physics running until 8pm Wednesday when
Machine studies for collider physics will start.